The total number of seats in a basketball sports arena is 12,000. The arena is divided into three sections, court sides, end zone and balcony, and there are twice as many balcony seats as court side seats. For the conference championship game, ticket price were $10.00 for court side, $8.00 for balcony and $7.00 for end-zone. If the arena was sold out for the game and the total receipts were $99,000 how many seats were court-side???.... thank you ^^
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

Courtside seats ----------x $10
balcony seats ------------2x$8
end zone seats ------------y $7
capacity 12000
Receipts =$99,000
..
x+2x+y = 12000
3x+y=12000
10x+8*2x+7y=99000
26x+7y= 99000
3x+y=12000 .............1
26x+7y= 99000 .............2
Eliminate y
multiply (1)by -7
Multiply (2) by 1
-21x-7y= -84000
26x+7y= 99000
Add the two equations
5x=15000
/5
x= 3,000
plug value of x in (1)
3x+y= 12000
9000+y=12000
y=12000 -9000
y=3000

Courtside seats ----------x ----3000
balcony seats ------------2x----6000
end zone seats ------------y----3000

CHECK

3 x + 1 y = 12000
9000 3000 = 12000

Answer by stanbon(75887)

The total number of seats in a basketball sports arena is 12,000. The arena is divided into three sections, court sides, end zone and balcony, and there are twice as many balcony seats as court side seats. For the conference championship game, ticket price were $10.00 for court side, $8.00 for balcony and $7.00 for end-zone. If the arena was sold out for the game and the total receipts were $99,000 how many seats were court-side???....
----
Seat Equation: c + e + b = 12000
Seat Equation:-2c+ 0 + b = 0
Value Equation:10c+7e+8b = 99,000
========================================
I used a matrix function on a TI-84 to get:
court = 3000 seats
end zone = 3000 seats
balcony = 6000 seats
===========================
Cheers,
Stan H.
=============